# Linying Lv

### v3

@@ 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
Analyst information value may vary in the cross-section of stocks. Prior literature suggests that information processing and dissemination costs vary with firm size (Zeghal, 1984), and analyst opinions may trigger larger market reactions for smaller stocks due to their relatively limited public information environment (Stickel, 1995). Figure A9 illustrates this relationship through a scatter plot of information value against stock market capitalization. While the plot reveals a positive association between market equity and information value, the relationship exhibits notable nonlinearities. To formally examine this relationship, I estimate regressions of information value on stock size, incorporating firm and year fixed effects and the book-to-market ratio control.
Table 10 shows that stocks with larger market equity have significantly higher information value from analysts. The economic magnitude is substantial: estimates from column (4) indicate that a 1% increase in market equity corresponds to a 0.864% increase in analyst information value. Columns (5)-(8) reveal that textual information value exhibits greater size sensitivity compared to the combined numerical and textual information value. To better understand the mechanisms driving this positive size elasticity, I decompose the information value into two components: explained return volatility and price impact. The decomposition can be expressed logarithmically as:
$\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (11)
Panel B of Table 10 shows that higher analyst information value in larger stocks stems primarily from their greater liquidity. While analyst reports do not better explain return volatility in large stocks, the reduced price impact of liquid stocks enhances their information value.
@@ 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
Prior literature suggests that analysts who make bold forecasts incorporate private information more comprehensively than those who herd (Clement and Tse, 2005). This section examines the relationship between information value and analysts’ herding behavior.
To analyze cross-sectional variation in analyst characteristics, I estimate regressions of stock-analyst-day information value on forecast boldness indicators, incorporating firm, analyst, and year fixed effects. Following Jacob et al. (1999), who suggests larger brokerage firms provide superior research, I include broker size as a control variable. The analyst-specific information value measure is defined as:
$\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}}.$ (12)
Table 11 presents the regression results. Panel A, columns (1)-(4), reveals significantly positive coefficient estimates on $Bold$, indicating more strategically valuable information when analysts provide bold forecasts. Quantitatively, information from bold reports demonstrates 29% higher value compared to herding reports—a magnitude that persists when examining textual content alone. The insignificant coefficients for $Brokersize$ suggest that broker size is related to analyst information value.
To understand the underlying mechanisms, Panel B decomposes the relationship between analyst boldness and information value into informativeness and price impact components. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts issue bold forecasts when possessing relevant private information. Columns (5)-(8) reveal no statistically significant relationship between boldness and price impact, suggesting that the enhanced value of bold forecasts stems primarily from their information content rather than market microstructure effects.
@@ 3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. Accordingly, this section evaluates the information value of analyst reports issued promptly following earnings announcements and beyond that.
Figure 5 examines different samples of analyst reports categorized by their timing relative to earnings announcements, from reports released in the first week after earnings announcements to those issued in the $13^{th}$ week. Reports issued in the first post-earnings week exhibit the highest information value, with a sample mean of \$0.84 million. Reports from subsequent weeks show a sharp decline in value. This pattern again suggests that markets place premium value on timely post-earnings analyst reports.
To formally test the observed pattern, I estimate regressions of log information value on an indicator variable $Week$, which identifies reports released within one week after earnings announcements. Following Banerjee (2011) and Beckmann et al. (2024), I include earnings announcement day trading volume ($TradingVolume$) as a proxy for investor disagreement. The model incorporates an interaction term ($Week\times TradingVolume$) to capture the joint effect of post-earnings timing and investor disagreement on analyst information value. The specification builds on Chen et al. (2010)’s finding that analysts’ interpretative role becomes more critical with increasing financial information complexity. The model includes stock and year fixed effects.
Table 12 presents regression results. Panel A documents significantly higher information value for reports issued within the first post-earnings week, as evidenced by positive and significant coefficients on $Week$ across all models. The positive and significant interactions between $Week$ and $TradingVolume$ indicate that information value increases with trading volume on earnings announcement days. The results are consistent with investors placing greater reliance on analyst interpretations during periods of heightened disagreement.
Panel B decomposes the earnings announcement effects into return variance and price impact, showing significant increases in both components. The coefficient magnitudes indicate that the timing effect is primarily driven by higher explained return volatility, rather than changes in price impact. These findings lead to conclude that investors value the most of analysts’ interpretation role during periods of heightened post-earnings uncertainty.
@@ 4 Conclusion
This paper examines both the quantitative and qualitative content of analyst reports. I find that report text contains more market-valued information content than quantifiable forecast revisions economically, especially shortly following earnings announcements. A Shapley value decomposition identifies analysts’ interpretation of income statements as the most valuable component of written reports.
I further translate the information content into economic terms. Using a measure of explained return volatility relative to price impact, I find that early acquisition of the analyst reports for an average S&P 100 constituent firm yields an annualized profit larger than \$6.89 million. Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets. The substantial potential profits from early access to analyst reports raise concerns about selective distribution practices—a direct violation of FINRA Rule 2241, which mandates research objectivity and prohibits preferential distribution to certain clients. This study presents a hypothetical scenario for quantifying information value rather than an assessment of market practices. By documenting the magnitude of potential profits from privileged access, my research emphasizes the importance of regulations for fair disclosure.
@@ A. Additional Figures
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
@@
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
@@
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
@@
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
Figure A2: Large Language Model Architecture
This figure shows an example of how large language models process a sentence from an analyst report on Nvidia. The model breaks this sentence into tokens, and then generates two types of embeddings: token embeddings and position embeddings. These embeddings then pass through the core of the model - a stack of transformer layers. The output of each transformer layer is a series of vectors representing the tokens in their full context.
@@
This figure shows the estimated information value of analyst reports for Fama-French 12 industries from 2015Q1 to 2023Q4, quantified in millions of dollars. The solid green line marks the sample mean value of information, with green bars showing the 95% confidence intervals. The horizontal dashed red line marks the sample mean of stocks from all industries. Information value estimates are derived using the delta method and adjusted to 2020 levels using the Consumer Price Index (CPI).
Figure A4: Shapley Values Scaled by Length
This figure presents the SHAP values of topics scaled by the total number of sentences (Panel A) and the total number of tokens (Panel B) for topics with more than 100,000 sentences. The scaling process involves two steps: First, normalize the Shapley value of each topic by dividing it by the number of sentences in that topic. Then, adjust the normalized values to ensure the total sum of SHAP values remains consistent. This ensures that the scaled values reflect the relative importance of topics while preserving the magnitude of total $R^{2}_{\mathrm{OOS}}$.
Figure A5: Shapley Values for Topic Importance across Industries
This figure shows the relative importance of 17 topics in analyst reports across the Fama-French 12 industries, quantified using the SHAP value. For each industry, the bars represent the SHAP values of different topics, with their sum equaling the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings. The red line shows each topic’s SHAP value as a proportion of the total $R^{2}_{\mathrm{OOS}}$.
Figure A6: $R^{2}_{\mathrm{OOS}}$ of Sentence-Segmented Embeddings by Year
This figure shows the $R^{2}_{\mathrm{OOS}}$ of sentence-segmented report embeddings from 2015 to 2023. The color gradient indicates the magnitude of $R^{2}_{\mathrm{OOS}}$. The red line represents the SHAP value of the Income Statements Analyses topic across the years. The shaded area marks the pandemic recession in 2020.
Figure A7: Shapley Values for Topic Importance by Analyst Characteristics
@@
This figure illustrates how 17 different topics contribute to the predictive power of analyst reports across four subsamples, measured using SHAP values. The subsamples compare analysts with above-median (Panel A) versus below-median work experience (Panel B), and large (Panel C) versus small brokerage firms (Panel D), with subsamples determined by yearly median. For each subsample, the bars show the SHAP values of individual topics, whose sum equals the out-of-sample R-squared ($R^{2}_{\mathrm{OOS}}$) from the sentence-segmented report embeddings. The red line indicates each topic’s relative importance, calculated as the topic’s SHAP value divided by the total $R^{2}_{\mathrm{OOS}}$.
@@
Figure A8: Alternative Analyst Information Value Estimations Over Time
@@
This figure presents the alternative estimations of analyst information value from 2015Q1 to 2024Q4, reported in millions of dollars and adjusted to 2020 dollars. Quarterly mean approximations are calculated using the delta method. TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Figure A9: Analyst Information Value by Stock
@@
This figure shows the estimated information value of analyst reports for individual stocks from 2015Q1 to 2023Q4, quantified in millions of dollars. Estimates are derived using the delta method. Market equity, averaged daily at the time of each report release, is reported in billions of dollars. All values are adjusted to 2020 levels using the Consumer Price Index (CPI).

### v4

@@ 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
The economic value of analyst information may vary across the cross-section of stocks. Prior literature offers competing hypotheses regarding firm size. On the one hand, information acquisition and analysis are typically more costly for smaller, less-followed firms, suggesting analysts might focus on larger firms (Zeghal, 1984). On the other hand, the relative scarcity of public information for smaller firms could make analyst opinions more influential and thus more valuable to the market (Stickel, 1995). This creates an empirical question as to whether information value is positively or negatively related to firm size.
Figure A9 illustrates this relationship through a scatter plot of information value against stock market capitalization. While the plot reveals a positive association between market equity and information value, the relationship exhibits notable nonlinearities. To formally test this relationship, I estimate regressions of the information value measure on firm size (log market equity), and controlling for the book-to-market ratio, firm, and year fixed effects.
The results in Table 9 confirm a strong positive relationship. The coefficient estimates in Column (4) imply an elasticity of 0.864, meaning a 1% increase in market equity is associated with a 0.864% increase in analyst information value. Notably, Columns (5) through (8) show that this elasticity is even stronger for text information value, suggesting that the interpretive role of analysts is particularly important for larger, often more complex firms.
To better understand the mechanisms behind this relationship, I decompose information value into two components: the explanatory power of the report and the price impact of trading. Expressed in logarithmic form, the decomposition is given by:
$\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\hat{r}_{it}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (10)
This allows for testing whether the higher information value for large firms stems from an information channel (analysts providing better or more impactful information) or a liquidity channel (trading on the information is cheaper due to lower price impact).
Panel B of Table 9 shows that the higher information value for large-cap stocks arises mainly from lower price impact, not greater explanatory power. That is, while analyst reports do not better explain return variance for larger firms, their improved liquidity reduces transaction costs, thereby enhancing the value of analyst information from a trading perspective.
Panel B of Table 9 presents results of the decomposition analysis. The findings are striking: the positive relationship between firm size and information value arises mainly from the liquidity channel. While analyst reports do not exhibit significantly greater explanatory power for larger firms, the price impact of trading is substantially lower for these high-liquidity stocks. This reduced transaction cost enhances the value of analyst information. As my sample is composed of large stocks, it is crucial to recognize that the economic value of analyst information likely decreases for smaller, less liquid firms.
@@ 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
This section explores how the economic value of analyst reports varies with analyst characteristics, focusing on bold and herding forecast behavior. Prior research suggests that bold forecasts are more likely to be based on an analyst’s private information, whereas herding behavior often reflects a reliance on public signals or consensus views (Clement and Tse, 2005). I therefore test the hypothesis that bold reports generate greater strategic information value.
To examine this cross-sectional variation, I regress the stock-analyst-day level information value on an indicator for forecast boldness, controlling for brokerage size as well as firm, analyst, and year fixed effects. The analyst-specific information value measure is defined as:
$\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}},$ (11)
where $\hat{r}_{ijt}$ is the out-of-sample CAR estimation for report $j$ of stock $i$ on day $t$.
The regression results are presented in Table 10. Panel A shows that the coefficient on the boldness indicator (Bold) is positive and highly significant across all specifications. The estimated magnitude implies that, on average, bold analyst reports are 29% more valuable than their herding counterparts. This pattern holds when analyzing the value derived from textual content alone. In contrast, the coefficient on brokerage size is consistently insignificant, suggesting that after controlling for other factors, the size of brokerage firms does not materially influence the economic value of analysts’ reports.
To understand the channel through which boldness creates value, Panel B decomposes the information value into its two components: explained return variance (the information channel) and price impact (the liquidity channel). The results clearly indicate that the effect is driven by the information channel. Columns (1)–(4) show that bold reports are associated with significantly larger explained return variance, reinforcing the view that they contain meaningful signals. Conversely, Columns (5)–(8) reveal no systematic relationship between boldness and price impact, going against the possibility that the higher value of bold reports stems from differential market liquidity or trading frictions.
@@ 3.3.3 The Value of Analyst Information around Earnings Announcements
Building on earlier findings that the statistical information content of analyst reports peaks shortly after earnings announcements, this section examines whether the economic value of these reports follows a similar pattern. I investigate how the value of analyst information varies based on its timeliness relative to corporate earnings disclosures.
Figure 5 plots the average information value of analyst report subsamples across 13 weekly bins following earnings announcements. Reports issued in the first week show the highest average value, at approximately \$0.84 million. This value declines sharply in subsequent weeks, suggesting that the market places a significant premium on timely analyst reports that digest and interpret corporate news.
To formally test this observation, I estimate panel regressions of log information value on a post-earnings week indicator ($Week$), which equals one if the report is issued within one week of the earnings announcement. Following prior literature that suggests analyst interpretation is most valuable amidst uncertainty (Chen et al., 2010), I also test if this effect is amplified during periods of high investor disagreement. I proxy for disagreement using trading volume on the earnings announcement day ($TradingVolume$) and include an interaction term ($Week\times TradingVolume$), following Banerjee (2011) and Beckmann et al. (2024). All specifications include stock and year fixed effects.
The regression results are presented in Table 11. Panel A documents significantly higher information value for reports issued within the first post-earnings week, as evidenced by positive and significant coefficients on $Week$ across all models. Furthermore, the positive and significant coefficient on the interaction term indicates that this timeliness premium is even greater when there is high investor disagreement surrounding earnings announcements. This is consistent with investors placing the greatest reliance on analyst sense-making during periods of heightened uncertainty.
Panel B decomposes this finding to identify the channel through which timeliness creates value. By regressing the two components of the information value measure on the same set of variables, I can distinguish between an information channel and a liquidity channel. The results show that the timeliness effect is driven primarily by the information channel: reports issued shortly in the post-earnings announcement window are associated with a significant increase in explained return variance, while changes in price impact are relatively modest.
@@ 4 Conclusion
This paper examines both the quantitative and qualitative information content of analyst reports. I find that the textual content of analyst reports carries more economically significant, market-valued information than quantifiable forecast revisions. This effect is especially pronounced in the period immediately following earnings announcements. A Shapley value decomposition identifies analysts’ interpretation of income statements as the most valuable topic within written reports.
To translate these findings into economic terms, I adopt a measure of information value accounting for both the explained return variance and the price impact. The results suggest that early acquisition of analyst reports for an average S&P 100 constituent stock yields annualized profits exceeding \$6.89 million. The value of analyst information is higher for large-cap firms, bold analyst forecasts, and when investors’ disagreement is high.
These findings have important implications for regulators. By quantifying the substantial economic value of privileged access to analyst reports, I reveal strong incentives for selective distribution, a practice prohibited by FINRA Rule 2241. It is important to note that the analysis provides a hypothetical framework for understanding the potential value of such information, not an empirical assessment of market practices. Nevertheless, the magnitude of these potential profits reinforces the fundamental principle that equal access to value-relevant information is essential for market integrity.
@@ A. Additional Figures
Figure A1: Large Language Model Architecture
This figure demonstrates how a large language model generates context-aware representations using a sample sentence from an analyst report on NVIDIA. The model first tokenizes the sentence and generates two types of embeddings: token embeddings and positional embeddings. These are then passed through a stack of transformer layers, the core of the architecture. The output of each transformer layer is a sequence of context-aware vectors, one for each token, capturing its meaning concerning the full context.
@@
Figure A2: Word Clouds of Analyst Report Topics
This figure displays word clouds for 16 key topics commonly discussed in analyst reports. Each word cloud highlights the most frequently occurring terms within each topic, with word size proportional to term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
@@
Figure A2: Word Clouds of Analyst Report Topics
This figure displays word clouds for 16 key topics commonly discussed in analyst reports. Each word cloud highlights the most frequently occurring terms within each topic, with word size proportional to term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
@@
Figure A2: Word Clouds of Analyst Report Topics
This figure displays word clouds for 16 key topics commonly discussed in analyst reports. Each word cloud highlights the most frequently occurring terms within each topic, with word size proportional to term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
@@
Figure A2: Word Clouds of Analyst Report Topics
This figure displays word clouds for 16 key topics commonly discussed in analyst reports. Each word cloud highlights the most frequently occurring terms within each topic, with word size proportional to term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
@@
This figure shows the estimated information value of analyst reports across Fama-French 12 industries from 2015Q1 to 2023Q4, with values reported in millions of 2020 dollars and estimated via the delta method. The solid green line indicates the mean information value for each industry, while the green bars represent 95% confidence intervals. The horizontal dashed red line denotes the cross-industry average.
Figure A4: Shapley Values Scaled by Topic Length
This figure adjusts the topic Shapley values to measure their per-sentence or per-token contribution rather than their aggregate importance, focusing on topics with over 100,000 sentences. Panel A presents the values scaled by the number of sentences in each topic, while Panel B presents them scaled by the number of tokens. This two-step scaling process involves first normalizing each topic’s total Shapley value by its length (sentences or tokens) and then rescaling the results to maintain the original sum of all values, thereby preserving the model’s total explanatory power.
Figure A5: Topic Importance via Shapley Value across Industries
This figure decomposes the information content ($R^{2}_{\mathrm{OOS}}$) of analyst reports into contributions from 17 topics for each of the Fama-French 12 industries. Within each industry’s panel, the height of a bar represents the Shapley value for a specific topic, indicating its contribution to the total out-of-sample $R^{2}$ for reports in that sector. The red line shows each topic’s SHAP value as a proportion of the aggregated $R^{2}_{\mathrm{OOS}}$.
Figure A6: Information Content of Sentence-Segmented Embeddings by Year
This figure shows the evolution of the total out-of-sample explanatory power ($R^{2}_{\mathrm{OOS}}$) of sentence-segmented analyst reports from 2015 to 2023. Specifically, the red line isolates the contribution of the ‘Income Statement Analyses’ topic, as measured by its annual Shapley value. The shaded area denotes the 2020 pandemic recession.
Figure A7: Topic Importance via Shapley Value by Analyst Characteristics
@@
This figure illustrates how 17 different topics contribute to the analyst report information content across six subsamples. The subsamples compare analysts with above-median (Panel A) versus below-median work experience (Panel B), large (Panel C) versus small brokerage firms (Panel D), and the primary valuation methodology used (Panel E: Multiple-Based vs. Panel F: DCF-Based). The splits for experience and brokerage size are based on the yearly median. For each group, bars represent the SHAP values of individual topics, with the red line indicating each topic’s relative importance scaled by total $R^{2}_{\mathrm{OOS}}$.
@@
Figure A8: Alternative Estimations of Analyst Information Value Over Time
@@
This figure presents several alternative estimations of the quarterly analyst information value from 2015Q1 to 2024Q4, serving as a robustness check to the primary measure. These alternatives modify key components of the main calculation. The ‘TAQ volatility’ series uses realized volatility from one-minute log returns ($\sigma_{v}^{2}$) to calculate explained return volatility as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$. The ‘EMO’ and ‘CLNV’ series use alternative trade-signing algorithms for the price impact calculation developed by Ellis et al. (2000) and Lee and Ready (1991), respectively. All estimates are reported in millions of 2020 dollars, with quarterly means approximated via the delta method.
Figure A9: Analyst Information Value and Stock Market Equity
@@
This figure plots the estimated information value of analyst reports against the market equity for each individual stock, covering the period from 2015Q1 to 2023Q4. The information value (vertical axis) for each stock is estimated using the delta method and reported in millions of 2020 dollars, adjusted using the CPI. The market equity (horizontal axis) is the firm’s average daily market capitalization at the time of each report release, reported in billions of dollars.